INVESTIGATIONS OF THE COLLISION-INDUCED ABSORPTION OF O_{2} NEAR 6.4 μm IN PURE O_{2} AND O_{2}/N_{2} MIXTURES

C.L. LUGEZ, A.M. SOLODOV, \underline{W.J. LAFFERTY}, AND G.T. FRASER, OPTICAL TECHNOL-OGY DIVISION, NATIONAL INSTITUTE OF STANDARDS AND TECHNOLOGY, GAITHERS-BURG, MD 20899.

Measurements of the collision-induced absorption of O_{2} in the vicinity of the electric-dipole-forbidden vibrational fundamental have been made in pure O_{2}, O_{2}/N_{2} and Ar/O_{2} mixtures. The goal is to provide reliable absorption coefficients to allow accurate retrievals of stratospheric concentration profiles of H_{2}O vapor, NO_{2}, and sulfate aerosols from limbviewing filter radiometers, such as those mounted on the UARS and future EOS platforms. Here, spectra were recorded at a resolution of 0.5 cm^{−1} using a Fourier-transform infrared spectrometer. Measurements were made using an optical pathlength of 84 m, sample densities up to approximately 8 amagats, and sample temperatures between 225 K and 300 K. The present results are in good agreement with the earlier less precise measurements of Orlando et al. on O_{2} and O_{2}/N_{2} mixtures and Thibault et al. on pure O_{2}. Studies of Ar/O_{2} mixtures indicate that the observed structure on the high-frequency shoulder of the band, previously attributed to dimer bound states by Henderson and Ewing, is more likely due to line mixing of pure quadrupole transitions of O_{2}, as originally suggested by Vigasin.